twice a number minus a second number is -1. twice the second number added to three times the first number is 9. Find the numbers.
Answer by fractalier(6550)

Call the numbers x and y. Then we have
2x - y = -1 and
2y + 3x = 9
Let's rearrange the first and get
y = 2x + 1
and substitute that into the second equation...we get
2(2x+1) + 3x = 9
4x + 2 + 3x = 9
7x + 2 = 9
7x = 7
x = 1
so that
y = 2(1) + 1 = 3